# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v1

@@ IV.1 Physics Models
Figure A.2 in the appendix shows AVFD simulation results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. Three values of $n_{5}/s$ are simulated. The results are consistent with those in Ref. [30]. The intercept at $v_{2,{\rm single}}=0$ is sensitive to CME with $v_{2}$ induced background significantly reduced. For $n_{5}/s=0$ the intercept is consistent with zero as one would expect if the background is fully removed. For $n_{5}/s=0.1$ and 0.2, the intercepts are found to be $1.34\times 10^{-4}$ and $5.70\times 10^{-4}$, respectively. The CME signal is expected to scale with $(n_{5}/s)^{2}$. The ratio of the ESS intercept for $n_{5}/s=0.2$ to that for $n_{5}/s=0.1$ is $4.25\pm 0.13$, deviating from the expected ratio of 4 by about 2 standard deviations.
@@
Figure A.2 shows the AVFD results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ from a phase space separated from that of the POIs. Specifically, the $\hat{q}_{2}^{2}$ is calculated using particles from $|\eta|<0.3$, and the POIs to calculate $\Delta\gamma$ and $v_{2}$ are from $20.3<|\eta|<$ 2 (upper cut of 2 instead of typically 1 to increase statistics). The intercept for the $n_{5}/s=0$ case is consistent with zero as expected. The intercepts for $n_{5}/s=0.1$ and 0.2 are $(1.63\pm 0.19)\times 10^{-4}$ and $(6.61\pm 0.19)\times 10^{-4}$, respectively. The statistical uncertainties are larger than those from ESS because of the larger range of projection to the intercept. The intercept ratio between $n_{5}/s=0.2$ and 0.1 is $4.06\pm 0.49$, consistent with the expected ratio of 4. It is interesting to note that the fit slopes from both ESS and ESE vary with $n_{5}/s$ significantly, whereas it is expected to measure the background strength. This is likely because the CME signal is also affecting $v_{2}$ of particles in the final state [60] and thus affects the slope as well.
@@
Figure A.4 shows the AMPT results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The intercepts are all negative except the 20–30% centrality at 27 GeV which is not statistically significant. Since AMPT does not have CME signals, the results suggest that the ESS intercept may not measure only CME signals but is also sensitive to backgrounds. The fit slope increases with decreasing centrality and is larger at 27 GeV than at 200 GeV; this is expected because the $\Delta\gamma$ background is approximately inversely proportional to multiplicity.
@@
Figure A.4 shows the AMPT results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ (calculated from particles within $|\eta|<0.3$) from a phase space separated from that of the POIs (taken from $10.3<|\eta|<$ 1). The intercepts are consistent with zero as expected, albeit large statistical uncertainties. The slope fit errors are too large to draw firm conclusions. We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward rapidities $43<|\eta|<$ 4, keeping the same POI acceptance. The results are similar but with larger errors.
@@
Figure A.5 shows the EPOS results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The POI and $\hat{q}_{2,{\rm pair}}^{2}$ pseudorapidity acceptance is taken to be $|\eta|<2$ to increase statistics. The intercepts appear to be positive though with large uncertainties, opposite to those in AMPT. This again suggests that the ESS intercept may be sensitive to backgrounds since EPOS does not have CME signals. There is an indication of an increasing fit slope with decreasing centrality, however, the uncertainties are large.
@@
Figure A.6 shows the EPOS results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ from a phase space separated from that of the POIs. The POI pseudorapidity acceptance is taken to be $20.3<|\eta|<$ 2 to increase statistics. The intercepts are consistent with zero as expected, albeit large statistical uncertainties. The slope fit errors are too large to draw firm conclusions.
@@
Figure A.8 shows the HYDJET++ results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The intercepts are consistent with zero at 200 GeV but are all positive at 27 GeV. Since HYDJET does not have CME signals, the results again suggest that the ESS intercept may not measure only CME signals but is also sensitive to backgrounds. The HYDJET results further suggest that the degree of the sensitivity of the intercept to backgrounds depends on the collision energy, i.e. details of the event content. The fit slope increases with decreasing centrality and is larger at 27 GeV than at 200 GeV, as expected, similar to the AMPT results.
@@
Figure A.8 shows the HYDJET results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ (calculated from particles within $|\eta|<0.3$) from a phase space separated from that of the POIs (taken from $10.3<|\eta|<$ 1). The intercepts are all consistent with zero as expected, except the 40–50% centrality at 27 GeV where the intercept is displaced from zero but only with less than 2 standard deviations. The fit slope increases with decreasing centrality and is larger at 27 GeV than at 200 GeV, as expected.
@@
We have also used the three-particle correlator method to calculated $\Delta\gamma$ and the two-particle cumulant to calculate $v_{2}\{2\}$, without relying on the known EP at $\psi=0$. The results are shown in Fig. A.10. The results are consistent with those in Fig. A.8 calculated with the known $\psi=0$, with somewhat larger statistical uncertainties.
@@
We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities $43<|\eta|<$ 4, keeping the same POI acceptance. The results are shown in Fig. A.10. The intercepts are all consistent with zero, except the 30–40% centrality bin at 27 GeV where the intercept is positive away from zero with 2 standard deviations.
@@ IV.2 Discussions on Physics Model Results
We summarize the intercepts and the intercepts divided by the overall $\Delta\gamma$ magnitude from the ESS method in Fig. 6 and from the ESE method in Fig. 6. We observe that the ESS intercepts $\Delta\gamma_{\textsc{ess}}$ are close to zero but not always consistent with zero. The ESE intercepts $\Delta\gamma_{\textsc{ese}}$ are mostly limited by large statistical uncertainties.
Figure 5: ESS intercepts. Shown are intercepts (upper panels) and intercept normalized by the inclusive $\Delta\gamma$ magnitude from the ESS method for various physics models. For AVFD ($n_{5}/s=0)$, the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For AMPT, EPOS, and HYDJET, the results are from the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. The POI $\eta$ acceptance is $|\eta|<1$ for AVFD, AMPT, and HYDJET, and $|\eta|<2$ for EPOS, and the $p_{\rm T}$ acceptance is $20.2<p_{\rm T}<$ 2 GeV/$c$ for all models. The $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs.
@@
Figure 6: ESE intercepts. Shown are intercepts (upper panels) and intercept normalized by the inclusive $\Delta\gamma$ magnitude from the ESE method for various physics models. For AVFD ($n_{5}/s=0$), the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For AMPT, EPOS, and HYDJET, the results are from the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. The POI $\eta$ acceptance is $10.3<|\eta|<$ 1 for AVFD, AMPT, and HYDJET, and $20.3<|\eta|<$ 2 for EPOS, and the $p_{\rm T}$ acceptance is $20.2<p_{\rm T}<$ 2 GeV/$c$ for all models. The $\hat{q}_{2}^{2}$ is calculated using particles from $|\eta|<0.3$ for all models (and also $43<|\eta|<$ 4 for AMPT and HYDJET) and $20.2<p_{\rm T}<$ 2 GeV/$c$. Several points are missing as they are outside the frame but consistent with zero with large uncertainties.
As discussed in the introduction, the ESS method uses the POIs to compute the event selection $\hat{q}_{2,{\rm pair}}^{2}$ variable. The $\hat{q}_{2,{\rm pair}}^{2}$ is correlated to the $\Delta\gamma$ and $v_{2,{\rm single}}$ of POIs and thus selects events based primarily on statistical fluctuations. The wide range of $v_{2,{\rm single}}$ and $\Delta\gamma$ are a result of statistical fluctuations by insisting on certain values of $\hat{q}_{2,{\rm pair}}^{2}$. Because of the statistical fluctuation nature, the $v_{2}$-induced backgrounds unnecessarily must be zero in the ESS intercept $\Delta\gamma_{\textsc{ess}}$; a model demonstration of this for the similar but simpler ES method can be found in Ref. [34]. It appears from the physics model studies that the value of $\Delta\gamma_{\textsc{ess}}$ can be negative, zero, or positive, depending on details of the model events. On the other hand, it can be proved that the flow-induced background must be zero at the ESE intercept $\Delta\gamma_{\textsc{ese}}$. One thus expects $\Delta\gamma_{\textsc{ese}}=0$ where nonflow is absent (which is the case for results using the known $\psi=0$). The ESE results are consistent with this expectation though the large statistical uncertainties prevent a firm conclusion.
One important difference among the models we studied is the various contributions from $K_{S}$. AVFD and AMPT keep the $K_{S}$ stable so there is no contamination from $K_{S}$ decay pions in the final state. EPOS and HYDJET, on the other hand, decay $K_{S}$ so there is maximum contamination of final-state pions (as we have used all pions in the model final state in our calculations). The intercepts from AMPT at both 200 and 27 GeV are negative, and the intercept from AVFD at 200 GeV is consistent with zero (AVFD is not ready for 27 GeV [61]). The intercepts from EPOS at 200 and 27 GeV are both positive. The intercept from HYDJET is consistent with zero at 200 GeV and positive at 27 GeV. In all models the strong-decay resonances (such as the $\rho^{0}$ resonance) are forced to decay. Relative particle abundances, for example, the $K_{S}$ over $\rho^{0}$ ratio, depends on energy; the lower the energy, the larger the low- to high-mass ratio. While very much dependent on models, it seems that $\Delta\gamma_{\textsc{ess}}$ increases with increasing $K_{S}/\rho^{0}$ ratio. It is thus conceivable, as resonances are a main background contribution to $\Delta\gamma$, that the relative resonance abundances could be important for the ESS intercept.

### v3

@@ IV.1 Physics Models
Figure A.2 in the appendix shows the avfd simulation results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The $\hat{q}_{2,{\rm pair}}^{2}$ is computed from the same POIs used for the $\Delta\gamma$ and $v_{2,{\rm single}}$ measurements. Three values of $n_{5}/s$ are simulated. The results are consistent with those of Ref. [36]. In avfd and all other physics models we studied, $\Delta\gamma$ appears to be linear as a function of $v_{2,{\rm single}}$. We fit the measured range of $0.20<v_{2,{\rm single}}<$ 0.2 to obtain the intercept $\Delta\gamma_{\textsc{ess}}$ at $v_{2,{\rm single}}=0$. The intercept is sensitive to CME with the $v_{2}$-induced background significantly reduced. For $n_{5}/s=0$, the intercept is consistent with zero as one would expect if the background was completely removed. For $n_{5}/s=0.1$ and 0.2, the intercepts are found to be $1.34\times 10^{-4}$ and $5.70\times 10^{-4}$, respectively. The CME signal is expected to scale with $(n_{5}/s)^{2}$. The ratio of the ESS intercept for $n_{5}/s=0.2$ to that for $n_{5}/s=0.1$ is $4.25\pm 0.13$, deviating from the expected ratio of 4 by about 2 standard deviations.
@@
Figure A.2 shows the avfd results analyzed using the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$. The $\hat{q}_{2}^{2}$ is calculated using particles from $|\eta|<0.3$, separate from the POIs from $20.3<|\eta|<$ 2 (upper cut of 2 instead of typically 1 to increase statistics) that are used to calculate $\Delta\gamma$ and $v_{2}$. The range of dynamical variation of $v_{2}$ is significantly smaller than the range of variation in $v_{2,{\rm single}}$ of the ESS method, which is predominantly statistical. The intercept $\Delta\gamma_{\textsc{ese}}$ is obtained from a linear fit to all data points. The intercept for the $n_{5}/s=0$ case is consistent with zero, as expected. The intercepts for $n_{5}/s=0.1$ and 0.2 are $(1.63\pm 0.19)\times 10^{-4}$ and $(6.61\pm 0.19)\times 10^{-4}$, respectively. The statistical uncertainties are larger than those from ESS because of the larger range of projection to the intercept. The intercept ratio between $n_{5}/s=0.2$ and 0.1 is $4.06\pm 0.49$, consistent with the expected ratio of 4. It is interesting to note that the fit slopes from both ESS and ESE vary with $n_{5}/s$ significantly, whereas it is expected to measure the background strength. This is likely because the CME signal particles have altered the strength of background-contributing correlations in the final state [69] and thus affected the slope as well.
@@
Figure A.3 shows the ampt  results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The intercepts are all negative; Negative intercepts were also found by the authors of Ref. [36, 71]. Since ampt does not have CME signals, the results suggest that the ESS intercept is sensitive not only to CME signals but also to backgrounds. The fit slope increases with decreasing centrality and is larger at 27 GeV than at 200 GeV; this is expected because the $\Delta\gamma$ background is approximately inversely proportional to multiplicity.
@@
Figure A.5 shows the ampt results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ calculated from particles within $|\eta|<0.3$ whereas POIs are taken from $10.3<|\eta|<$ 1. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions. We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities $53<|\eta|<$ 5 and $32<|\eta|<$ 3 for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.5 and are found to be similar to those with midrapidity $\hat{q}_{2}^{2}$ but with larger uncertainties.
@@
Figure A.7 shows the epos4 results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The POI and $\hat{q}_{2,{\rm pair}}^{2}$ pseudorapidity acceptance is taken to be $|\eta|<2$ to increase statistics. The intercepts appear to be positive, opposite to those in ampt. This again suggests that the ESS intercept may be sensitive to backgrounds since epos4 does not have CME signals. There is an indication of an increasing fit slope with decreasing centrality; however, the uncertainties are large.
@@
Figure A.7 shows the epos4 results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ from $|\eta|<0.3$. The POI pseudorapidity acceptance is taken to be $20.3<|\eta|<$ 2 to increase statistics. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions.
@@
Figure A.9 shows the hydjet++ results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in three centralities of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and 27 GeV. The intercepts are consistent with zero at 200 GeV but are all positive at 27 GeV. Since hydjet++ does not have CME signals, the results again indicate that the ESS intercept is sensitive not only to CME signals but also to backgrounds. The hydjet++ results further suggest that the degree of the sensitivity to backgrounds depends on the collision energy, or details of the event content. The fit slope increases with decreasing centrality and is larger at 27 GeV than at 200 GeV, as expected, similar to the ampt results.
@@
Figure A.9 shows the hydjet++ results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$. The intercepts are all consistent with zero, as expected, except the 20–30% and 40–50% centralities at 27 GeV where the intercept is displaced from zero but only with less than 2 standard deviations. The fit slope increases with decreasing centrality.
@@
We have also used the three-particle correlator method to calculate $\Delta\gamma$ and the two-particle cumulant to calculate $v_{2}\{2\}$, without relying on the known EP at $\psi=0$. The results are thus subject to nonflow effects. The results are shown in Fig. A.11 corresponding to the same hydjet++ data as in Fig. A.9. The results are consistent with those in Fig. A.9 calculated with the known $\psi=0$, with somewhat larger statistical uncertainties. This suggests that the nonflow effects are relatively small in hydjet++ within the statistical uncertainties of the current study.
@@
We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities of $53<|\eta|<$ 5 and $32<|\eta|<$ 3 for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.11 corresponding to the same hydjet++ data as in Fig. A.9. The intercepts are all consistent with zero, except the 40–50% centrality bin at 27 GeV where the intercept is positive but within 2 standard deviations from zero.
@@ IV.2 Discussions on Physics Model Results
Figure 5: ESE intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESE method for various physics models. For avfd ($n_{5}/s=0$), the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are from the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. The POI $\eta$ acceptance is $10.3<|\eta|<$ 1 for avfd, ampt, and hydjet++, and $20.3<|\eta|<$ 2 for epos4, and the $p_{\rm T}$ acceptance is $20.2<p_{\rm T}<$ 2 GeV/$c$ for all models. The $q_{2}^{2}$ is calculated using particles from $20.2<p_{\rm T}<$ 2 GeV/$c$ and $|\eta|<0.3$ for all models (and also forward/backward rapidities for ampt and hydjet++: $53<|\eta|<$ 5 at 200 GeV and $32<|\eta|<$ 3 at 27 GeV). Several points are missing as they are outside the frame but consistent with zero with large uncertainties.
Figure 6: ESS intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESS method for various physics models. For avfd ($n_{5}/s=0)$, the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are from the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. The POI $\eta$ acceptance is $|\eta|<1$ for avfd, ampt, and hydjet++, and $|\eta|<2$ for epos4, and the $p_{\rm T}$ acceptance is $20.2<p_{\rm T}<$ 2 GeV/$c$ for all models. The $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs.
We summarize the intercepts $\left\langle\Delta\gamma_{\textsc{ese}}\right\rangle$ and the intercepts divided by the overall $\Delta\gamma$ magnitude $\left\langle\Delta\gamma_{\textsc{ese}}/\Delta\gamma\right\rangle$ averaged over the 20–50% centrality range from the physics model studies by the ESE method in Fig. 5. The ESE intercepts are mostly limited by large statistical uncertainties. The large statistical uncertainties arise from the narrow dynamical range of $\left\langle v_{2}\right\rangle$ based on the $\hat{q}_{2}$ event selection in a given narrow centrality bin (see, particularly, Fig. A.5 lower panels, Fig. A.5, and Fig. A.7) and from the long extrapolation lever arm to the $\left\langle v_{2}\right\rangle=0$ intercept. This is particularly true for 27 GeV where some of the data points fall outside the graph frame of Fig. 5. The avfd, ampt, and epos4 models are computationally expensive, and it is difficult to accumulate statistics for a qualitative improvement beyond the present precision.
As discussed in Sect. II, the flow-induced background must be zero at the ESE intercept $\Delta\gamma_{\textsc{ese}}$ in the absence of nonflow. In most of our model studies and for all the model studies presented in Fig. 5, we used the known reaction plane $\psi=0$ in computing the $v_{2}$ and $\Delta\gamma$ variables, so nonflow is absent from these observables. While the statistical uncertainties are large, the ampt and hydjet++ $\Delta\gamma_{\textsc{ese}}$ using $\hat{q}_{2}^{2}$ from the $|\eta|<0.3$ range seem to be nonzero. The reason is likely due to nonflow effects caused by the ESE event selection using the midrapidity $\hat{q}_{2}^{2}$: because the $\eta$ distance is small, there can be nonflow correlation between a POI and a particle used in the $\hat{q}_{2}^{2}$ calculation. Such a correlation could bias the $\left\langle v_{2,{\rm res}}\right\rangle$ values of resonances contributing to CME backgrounds to be not strictly proportional to the final-state particle $\left\langle v_{2}\right\rangle$, resulting in a nonzero $\Delta\gamma_{\textsc{ese}}$ intercept. To reduce this effect, we also calculate $\hat{q}_{2}^{2}$ using particles at forward/backward pseudorapidities in ampt and hydjet++, the results of which are shown by the red points in Fig. 5. These intercepts $\Delta\gamma_{\textsc{ese}}$ are consistent with zero, confirming our hypothesis, though the uncertainty is large for the ampt result at 27 GeV.
Figure 6 summarizes the physics model results from the ESS method on the intercepts $\left\langle\Delta\gamma_{\textsc{ess}}\right\rangle$ and the intercepts divided by the overall $\Delta\gamma$ magnitude $\left\langle\Delta\gamma_{\textsc{ess}}/\Delta\gamma\right\rangle$, averaged over the 20–50% centrality range. As discussed in Sect. II, the ESS method uses the POIs to compute the event selection $\hat{q}_{2,{\rm pair}}^{2}$ variable, which are automatically (self-)correlated with the $\Delta\gamma$ and $v_{2,{\rm single}}$ of POIs. Thus, the $\Delta\gamma$ and $v_{2,{\rm single}}$ are affected by statistical fluctuations in $\hat{q}_{2,{\rm pair}}^{2}$ that is used to select events. The wide range of $v_{2,{\rm single}}$ and $\Delta\gamma$ are a result of these statistical fluctuations by insisting on certain values of the statistically fluctuating $\hat{q}_{2,{\rm pair}}^{2}$. The $v_{2,{\rm single}}$ value can go to zero as shown in Figs. A.2, A.3, A.7, and A.9. Due to the long projection lever arm, the statistical precision on the intercept $\Delta\gamma_{\textsc{ess}}$ is good.
We observe in Fig. 6 that the ESS intercepts are close to zero but not always consistent with zero, although the background is suppressed by 80% or more as shown in the lower panels. Because of the statistical fluctuation nature, the $v_{2}$-induced backgrounds unnecessarily must be zero in the ESS intercept $\Delta\gamma_{\textsc{ess}}$. A model demonstration of this for the similar but simpler ES method (see Sect. II) can be found in Ref. [44]. There, the reason is clear, as elucidated in Sect. II and by the middle-panel cartoon of Fig. 1: the resonance (generally background contributing source) $v_{2}$ values are proportional to the overall $v_{2}$ of final-state particles, and when events are binned in the observed, statistically fluctuating $v_{2}^{\rm obs}$ of final-state particles, the average $\left\langle v_{2}\right\rangle$ values of the background sources are positive at $v_{2}^{\rm obs}=0$ [44]. The situation is much more complicated here in the ESS method where $\Delta\gamma$ is plotted against $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}$. As discussed in Sect. III.7, the $\hat{q}_{2,{\rm pair}}^{2}$ calculated by Eq. 9 is composed of several components: (i) a flow component from hydrodynamic particle pairs, which is related to the single particle flow in a non-trivial way [69], (ii) a flow component of the resonances (CME background contributing sources), (iii) a nonflow component between the two pairs used to calculate the two-pair cumulant by Eq. 9 that contain daughters from, for example, the same resonance decay, and (iv) a self-correlation component from the sharing of a common particle between the two pairs. Because of these complicated nature in the event-by-event $\hat{q}_{2,{\rm pair}}^{2}$ variable, it is unclear how the $v_{2,{\rm single}}$ and $\Delta\gamma$ would behave, individually and intertwiningly, as functions of $\hat{q}_{2,{\rm pair}}^{2}$. It would not be expected a priori that the $\Delta\gamma$ as a function of $v_{2,{\rm single}}$, binned in $\hat{q}_{2,{\rm pair}}^{2}$, would project to a zero intercept in the absence of CME.
It appears from the physics model studies shown in Fig. 6 (where CME is absent) that the value of $\Delta\gamma_{\textsc{ess}}$ can, indeed, be negative, zero, or positive, presumably dependent of model details (model descriptions can be found in Sects. III.1III.4). Note the results in Fig. 6 are averaged from 20–50% centralities. One obvious difference in the simulated events, even with the same model, is the event centrality or multiplicity. As discussed in Sect. III.7 and mentioned above, the $\hat{q}_{2,{\rm pair}}$ quantity contains a self-correlation component besides being sensitive to single-particle $v_{2}$, multi-particle cluster $v_{2}$, and nonflow. The self-correlations are caused by sharing of the same POI between two pairs of POIs, and thus its effect is inversely proportional to multiplicity. To gain insight, we show in Fig. 7 the ESS intercept in ampt and hydjet++ as a function of centrality in terms of the average POI multiplicity. The centralities are defined by the charged hadron multiplicity within $|\eta|<0.5$ to be 0–5%, 5–10%, 10–20%, $\cdots$ (10% size), 70–80%, and 80-100% of the total MB event sample. As seen from Fig. 7, the nonzero $\Delta\gamma_{\textsc{ess}}$ intercept becomes stronger in more peripheral collisions. Within the same model, the $\Delta\gamma_{\textsc{ess}}$ seems to follow the same trend in POI multiplicity. This explains why the hydjet++ $\Delta\gamma_{\textsc{ess}}$ in 20–50% centrality range is consistent with zero at 200 GeV and positive at 27 GeV. This may suggest that the root reason for the nonzero intercept $\Delta\gamma_{\textsc{ess}}$ is not collision energy or centrality, but POI multiplicity. The multiplicity dependence of the ESS intercept $\Delta\gamma_{\textsc{ess}}$ clearly shows that it is sensitive to background correlations. The different signs of $\Delta\gamma_{\textsc{ess}}$ between ampt (negative) and hydjet++ (positive) is presumably due to different physics implemented in these models.
Figure 7: ESS intercept multiplicity dependence. The ESS intercept $\Delta\gamma_{\textsc{ess}}$ as a function of the average POI multiplicity from peripheral 80–100% to central 0–5% Au+Au collisions at 200 and 27 GeV from ampt (upper panel) and hydjet++ (lower panel). The most peripheral point for AMPT 200 GeV is outside the frame.
To conclude this part of the discussion, our physics model results show that the ESS intercept is not a well-controlled observable. Although the background in $\Delta\gamma_{\textsc{ess}}$ is strongly suppressed, one does not know how much background still remains in $\Delta\gamma_{\textsc{ess}}$, or even its sign. One does not know quantitatively what the ESS intercept exactly measures, so the ESS method is practically not useful to search for the CME.